21. You are about to deposit $139 into one of the following saving accounts to be left on deposit for 25 years, each bank offers an account with a different interest rate and compounding period. Assuming you want to maximize your wealth, how much money would be in the bank account that offers the best effective rate of return after 25 years?

Bank A. 9.5 percent rate compounded semi annually

Bank B. 9.4 percent rate compounded monthly

Bank C. 9.3 percent rate compounded daily

bank D. 9.2 percent rate compounded continuously
Answer #1

Bank a gives 9.5% rate compounded semi annually, then the effective annual rate is =(1+9,5 1,)?-1 = 1.09725-1 20109725 Bane AIf the bank o gives 9.2% rate compounded continuously the effective annual rate is .. .. . 20.092 . e 96365l - 1.096365-1= 94